OpenStudy (anonymous):

one angle of parallelogram is 23 less than six times the measure of the angle next to it. the sum of the measures of the angles is 180. find the measure of the adjacent angles of the parallelgram. The smaller angle measures _____ the larger angle measures_____ Help pleaseeeee

jimthompson5910 (jim_thompson5910):

one angle is x the other angle is 6x - 23 the two angles add to 180, so x + 6x - 23 = 180 x + 6x = 180+23 7x = 203 I'll let you finish

jimthompson5910 (jim_thompson5910):

tell me what you get